Make New Entry,
Make Followup Entry
User name Ole Hansen
Log entry time 16:00:26 on May 11, 2011
Entry number 358951
keyword=Day Shift Summary
SL: Ole Hansen
TO: Nikos Sparveris (TO)
3rd: Dave Gaskell (3rd)
SRC production data taking on 4He and slanted carbon.
Beam off for 2 hours due to injector studies. Frequent 5-10-minuteish
due to accelerator trips and other mishaps. Everything up and running on the
Hall A end, however.
- 08:00 Shift starts with production data taking in progress on
20 cm 4He.
- 08:20 Beam off for injector tuning, requested by Hall C/Qweak.
- 10:45 Beam returns. Taking approx. 1 hour of data on slanted carbon
- 12:30 Switching back to 4He target.
- 16:00 Shift ends with production data taking in progress, run #4385
- 4377 production, 4He, 1M, continued from owl shift
- 4378 production, 4He, 236k
- 4379 production, slanted carbon, 296k
- 4380 production, slanted carbon, 516k
- 4381 production, 4He, 235k
- 4382 production, 4He, 798k
- 4383 production, 4He, 817k
- 4384 production, 4He, 92k
- 4385 production, 4He, in progress